# ATLAS: Automated Tree-based Language Analysis System for C and C++ Source Programs

### v2

@@ A.1. CFG and DFG Generation Success Rate
A.2.1. Input Source Code
@@
A.2.2. LLVM IR-Based CFG (Clang)
@@
A.2.3. ATLAS Statement-Level CFG
@@ A.3. Prerequisites
A.4. Scenario A: Single File Analysis — Control Flow Graph Generation
@@
A.4.1. Input Source Code
We use factorial.cpp, a simple factorial function with a loop and conditional return:
@@ A.4.1. Input Source Code
A.4.2. Command
Terminal command
@@ A.4.2. Command
A.4.3. Output Files Generated
@@
A.4.4. Generated Control Flow Graph
@@ A.4.5. Interpretation
A.5. Scenario B: Combined Multi-Code View Generation
@@
A.5.1. Input Source Code
@@
A.5.2. Command
Terminal command
@@ A.5.2. Command
A.5.3. Generated Combined CFG+DFG Graph
@@ A.5.4. Interpretation
A.6. Scenario C: Multi-File Project Analysis
@@
A.6.1. Input Project Structure
@@
A.6.2. Command
Terminal command
@@ A.6.2. Command
A.6.3. Output Files Generated
@@
A.6.4. Generated Combined Project CFG
@@
A.6.5. Multi-File Analysis Benefits

### v3

@@ Appendix A ATLAS: Tool Demonstration Walkthrough
A.1.1. Input Source Code
@@
A.1.2. LLVM IR-Based CFG (Clang)
@@
A.1.3. ATLAS Statement-Level CFG
@@ A.2. Prerequisites
A.3. Scenario A: Single File Analysis — Control Flow Graph Generation
@@
A.3.1. Input Source Code
We use factorial.cpp, a file containing a simple factorial function with a loop and conditional return:
@@ A.3.1. Input Source Code
A.3.2. Command
Run the following terminal command:
@@ A.3.2. Command
A.3.3. Output Files Generated
@@
A.3.4. Generated Control Flow Graph
@@ A.3.5. Interpretation
A.4. Scenario B: Combined Multi-Code View Generation
@@
A.4.1. Input Source Code
@@
A.4.2. Command
Run the following terminal command:
@@ A.4.2. Command
A.4.3. Generated Combined CFG+DFG Graph
@@ A.4.4. Interpretation
A.5. Scenario C: Multi-File Project Analysis
@@
A.5.1. Input Project Structure
@@
A.5.2. Command
Run the following terminal command:
@@ A.5.2. Command
A.5.3. Output Files Generated
@@
A.5.4. Generated Combined Project CFG
@@
A.5.5. Multi-File Analysis Benefits